OpenStudy (anonymous):

suppose walmart stock is selling at $59 a share and mattel stock is selling at $28 a share. amy waller has a maximum of $6000 to invest. she wishes to purchase four times as many shares of walmart as of mattel. only whole shares of stock can be purchasd. a. How many shares of each will she purchase? b. How much money will be left over?

OpenStudy (anonymous):

59w + 28m <= 6000 w = 4m 59(4m) + 28m <= 6000 236m + 28m <= 6000 264m <= 6000 m <= 22.727272 m = 22 w = 88 a. 22 shares of Mattel, 88 shares of Walmart 6000 - (22(28) + 88(59)) = 192 b. $192 left over